OpenStudy (anonymous):

Choose the correct simplification of the expression g^5 h^4 / g^2 h^3 = and that's written as g to the 5th power times h to the 4th power divided by g squared times h to the third power.

OpenStudy (diyadiya):

Note: \[\Large \frac{x^m}{x^n}= x^{m-n}\]

OpenStudy (anonymous):

Oh wow.

OpenStudy (diyadiya):

\[\Large \frac{g^5h^4}{g^2h^3}= ?\]

OpenStudy (anonymous):

Yes! (:

OpenStudy (diyadiya):

Ok what can that be ?

OpenStudy (anonymous):

erm.. Can you give me the code, so I can do this easier please.

OpenStudy (diyadiya):

Do for g & h separately \[\Large \frac{g^5h^4}{g^2h^3}= \frac{g^5}{g^2} \times \frac{h^4}{h^3}\]

OpenStudy (diyadiya):

Can you use this formula ? \[\Large \frac{x^m}{x^n}= x^{m-n}\]

OpenStudy (diyadiya):

so what can be \[\Large \frac{g^5}{g^2}= ...\]

OpenStudy (diyadiya):

@careless850 ??

OpenStudy (anonymous):

I'm here. My computer is so laggy on this site when trying to reply.

OpenStudy (diyadiya):

\[\Large \frac{g^5}{g^2}=.. \]

OpenStudy (anonymous):

g^3

OpenStudy (diyadiya):

Right!!! :) \[\Large \frac{h^4}{h^3}=..\]

OpenStudy (anonymous):

h^1

OpenStudy (anonymous):

or h..

OpenStudy (diyadiya):

SO \[\Large \frac{g^5h^4}{g^2h^3}=..\]

OpenStudy (anonymous):

g^3 h^1

OpenStudy (diyadiya):

Right!!! :) Good

OpenStudy (anonymous):

g3h

OpenStudy (diyadiya):

or just \( g^3h \)

OpenStudy (anonymous):

Yeah, that. How do you do that? MathJax?

OpenStudy (diyadiya):

Yeah ,You can use the equation button

OpenStudy (diyadiya):

the blue colored one :)

OpenStudy (anonymous):

Well, good. I have another question for you! :D

OpenStudy (diyadiya):

Ok post it !!

OpenStudy (anonymous):

um.